OpenStudy (anonymous):

using only algebra find a cubic function with the given zeros. 3,-4,6

OpenStudy (jamesj):

Consider the quadratic x^2 - x - 2 Its roots are 2 and -1. That means we can write it as x^2 - x - 2 = (x-2)(x+1) Conversely if I had asked you: "What is the quadratic with roots 2 and -1?" you could have written down immediately that it has a factorization: (x-2)(x+1) Multiplying that back out, you would find it is x^2 - x - 2 So now for you equation, you have the three roots, so you can write down immediately its factorization.

OpenStudy (anonymous):

start with \[(x-3)(x+4)(x-6)=0\] and then multiply out